# Quasinormal spectra of a wormhole family: overtone features and a parameter-controlled redshift

### v1

@@ III.2 The spectral method
$(M_{0}+\omega M_{1}+\omega^{2}M_{2}+...+\omega^{p}M_{p})\phi=0$ (29)
III.3 Other methods: merits and demerits
Apart from the methods discussed above, there are a plethora of analytic and numerical methods that can be employed to compute the quasinormal spectra. One can compare numerically obtained QNMs with exact results for the few effective potentials which are analytically solvableFerrari and Mashhoon (1984); Heidari and Hassanabadi (2023). In some cases, the exactly solvable potentials are used as approximations for the real potentials in black hole perturbations (for example, in the Schwarzschild, a matching may be done using the Poschl–Teller potential, as shown many years ago in Ferrari and Mashhoon (1984) ). The QNMs found following this route however do not match accurately when compared with results using other numerical methods. In particular, the $Re(\omega)$ for the overtones show a mismatch.
IV Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
The problem of computing quasi-normal modes for the Bronnikov–Ellis wormhole has been tackled in many previous works. The authors of Kim (2008); Blázquez-Salcedo et al. (2018b) and Dutta Roy et al. (2020); Dutta Roy (2022) have presented the fundamental modes for the original Bronnikov–Ellis ($n=2$ case) and the generalised Bronnikov–Ellis wormholes respectively.
The authors of Baibhav et al. (2023) have argued that the overtones having lower real parts compared to the fundamental mode (true for $n=2$), fit away the non-linearities of the RD signal. The same argument will not be true for $n>2$ geometries (where $Re(\omega)$ for overtones are higher in value than the fundamental mode). Any possible future detection of such RD signals may further cement the role of overtones.
As stated above, while computing the $\omega_{QNM}$, we have used a frequency-domain method (spectral method (SM)) in conjunction with a time-domain method (Prony fit). While the frequency-domain method is indispensable for extracting all the modes (up to a certain overtone number), the time-domain provides us with some valuable insight into the relative amplitudes of the overtones, thus pointing to the fact that sometimes even higher-order overtones (in our case: even modes) can have greater importance than the lower-order (in our case: odd modes) ones.
@@ IV Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
In the series of figures below, we demonstrate the effective potentials and QNMs for scalar and axial gravitational perturbations. Figs. 2-5 are for $n=2$, Figs. $6-9$ for $n=4$, Figs. 10-13 are for $n=6$ and Figs. 14-17 for $n=8$. In the following section, we discuss the results here and try to arrive at specific conclusions on the nature of the QNMs and their link with the different geometries.
@@
V Inferences drawn from the QNM spectra: emphasis on overtones
In Section IV, we presented the fundamental modes and overtones of the generalised, ultrastatic, Bronnikov–Ellis wormhole. Now, one may ask whether these QNMs can help distinguish different geometries (i.e. different $n$ values) of this wormhole family. To answer this query, we focus on the plots ($Re(\omega)$ vs $-Im(\omega)$) given above in the previous section and try to identify some general trends in the QNM spectra. In our analysis, we denote the mode numbers with $p$, where the mode with the smallest value of $|Im(\omega)|$ (the longest-lived mode) is the fundamental mode (see e.g. Giesler et al. (2019)) with $p=0$ and modes with higher $|Im(\omega)|$ are the overtones.
For both scalar and axial perturbations, in the $n=2$ case, the real parts of the $\omega_{QNM}$ decrease with increase in mode number ($p$), while for the geometries with higher $n$ values, the higher overtones correspond to an increase in $Re(\omega)$.
From the tables in Appendix A, we notice that for a fixed large value of the angular momentum parameter ($m$), the fundamental modes of higher $n$ geometries are very close in value. However, these wormhole spacetimes can also be distinguished from each other if we take the overtones into account. As we move to larger $n(>2)$ values, we observe that the rate of increase of $Re(\omega)$ with the mode number gets expedited. This is true for both scalar and axial perturbations (see Figs. 6, 6, 10, 10…, 18).
@@ V Inferences drawn from the QNM spectra: emphasis on overtones
Having identified the broader trends from the plots, we can try to explain them from the effective potential. The decrease of $Re(\omega)$ with overtone mode number is only observed in the $n=2$ case, which has a comparatively narrower and sharper single-peaked potential (See Figs. 4, 4). This trend is reversed for all even $n>2$ wormholes, which have a comparatively broader and flatter effective potential with two peaks for scalar perturbation (See Figs. 8, 12, 16) and one (for higher values of m)/three (for smaller values of m) peak(s) for axial perturbations (See Figs. 8, 12, 16). On careful observation, one can notice that as we go to higher values of $n$, the central region of the potential becomes flatter. This can be a possible explanation behind the enhanced rate of increase of the real parts for higher $n$ geometries.
Fixing the value of $n$, if we increase the angular momentum parameter ($m$), the height of the effective potential increases, which is related to increasing the real part of the QNMs. The double peak feature also becomes less prominent for the scalar perturbation case (See Figs. 8, 12, 16). This may explain the decrease of $|Im(\omega)|$ with $m$ in the quasi-normal spectra. For axial perturbations in $n>2$ geometries, the transition of the potential from triple barrier to step-like and then to a single barrier, with increasing m (See Figs. 8, 12, 16), may cause the initial increase and subsequent decrease of $|Im(\omega)|$ for the fundamental modess with m.
@@
Increase n: of $Re(\omega)$ with the mode of $Re(\omega)$ with the mode
@@
Increase p: $Re(\omega)$ does not change $Re(\omega)$ increases after a
@@
V.1 Approximate Analytic Fit
@@
As mentioned above, the change in the QNMs with changing $n$ values may give us some valuable insights. To this end, we have analysed the dependence of $Re(\omega)$ on overtone mode number ($p$) and the angular momentum parameter ($m$), keeping the geometry parameter $n$ fixed.
@@
Here, we observe that the real parts of the quasi-normal frequencies decrease with overtone mode $p$ but increase with angular momentum $m$.
@@
For geometries with higher $n$ $(>2)$, the real parts increase with $m$ as well as with overtone mode number $(p)$. We notice, with increasing $p$ $Re(\omega)$ increases faster for $n=8$ case as compared to say, the $n=4$ case. We have also fitted the imaginary part with the geometry parameter and the real part, for different overtone numbers ($p$), for scalar perturbations.
@@
Thus far, we have (i) fitted $Re(\omega_{QNM})$ with angular momentum and overtone mode number and (ii) the $-Im(\omega_{QNM})$ with geometry parameter $n$ and $Re(\omega_{QNM})$. Hence, for given values of $n$, $m$ and $p$, we can numerically estimate the value of $\omega_{QNM}$ (scalar perturbation). The fitting functions discussed so far are approximate. They do not return exact results but provide information about overall trends, which are useful while interpreting our results.
V.2 Detectable Modes
@@
[…]
Advanced LIGO is sensitive to the frequency range of $10\hskip 5.69046ptHz$ to $10\hskip 5.69046ptkHz$. For the generalised Bronnikov-Ellis geometries, we have calculated the QNFs, and all of them can be within the aLIGO frequency range when the throat radius is roughly between $50M_{\odot}$ and $1000M_{\odot}$. For now, we choose $b_{0}=200M_{\odot}$ and present the frequencies and damping times for a few cases, as an illustration. In Table II below we show the values for $n=2$ and $m=5$ (scalar perturbation).
@@
Table 2: Frequency and Damping time for $n=2$ and $m=5$. These values are calculated from the Prony fit (details in Appendix A).
@@
For $n=8$ and $m=5$ case (scalar):
@@
Table 3: Frequency and Damping time for $n=8$ and $m=5$. These values are calculated from the prony fit (details mentioned in Appendix A).
@@
In the ringdown waveforms, we have represented the actual time-domain (obtained using the procedures mentioned in III.1) wave as the blue points, the fundamental mode as green lines and the resultant wave of the fundamental and overtones ($\omega$ and amplitudes computed using the Prony fit) as the magenta lines (Figs. 23,25). For both cases, the magenta lines better fit the time domain than only the fundamental mode. However, the difference between the green and magenta lines is only noticeable for the first few cycles, as the overtones have higher imaginary parts and decay faster. While the inclusion of overtones makes minimal improvement in the fitting for the $n=2$ case, the effect is more significant for the $n=8$ case. This is because for the $n=8$ case, we could find more overtones, and these overtones have amplitudes comparable to the fundamental mode, whereas, in the $n=2$ case, the amplitudes of the overtones are significantly lower than the fundamental mode. For $n=4,6$ cases, we computed $2-3$ overtones and their corresponding amplitudes using the Prony fit, and similar to the $n=8$ case, we observe that inclusion of the overtones significantly improves the fit at early times. For all the geometries (i.e. different $n$), the more overtones we could get (up to $p=4$), the better we were able to model the time-domain signal, indicating that obtaining more overtones may result in a better fit of the time-domain profile. However, we need to be cautious not to overfit our results by incorporating an excessive number of overtones.
One would expect the amplitudes of the modes to depend on our initial value of $\psi$ but not on which part of the time domain we are fitting. However, for the generalised Bronnikov–Ellis wormhole, only the amplitudes associated with the fundamental modes and 2nd overtones (to some extent) are time-independent. For almost all the other $\omega_{QNM}$, if we change the initial time of the fitting, the corresponding amplitudes remain stable up to some fitting domain, but beyond it, they begin to vary.
@@
Thus, we again fit the ringdown profiles with the modes with a stable amplitude
@@ VI Redshift
In the previous sections, we have computed the quasinormal modes for the massless generalised Bronnikov–Ellis wormhole. However, if the geometry under study is surrounded by a dark matter halo or any generic matter distribution, the QNMs calculated for the isolated compact object may differ from the observed values, as the environment may affect the generation of GW as well as their propagation. Such environments cause the spectrum to be red-shifted as documented in Barausse et al. (2014); Pezzella et al. (2025); Liu et al. (2025) for the black hole spacetime embedded in a halo.
@@
In the case of the Bronnikov–Ellis wormhole, similar redshifts in the fundamental mode can be found, if one considers the wormhole with non-zero ADM mass, as noted by the authors of Blázquez-Salcedo et al. (2018b). We will first find the QNMs (fundamental and overtones) for the $M\neq 0$ case in the $n=2$ Bronnikov–Ellis geometry. Thereafter, in our second example, we will find the QNMs for a modified $n=2$ geometry with a new redshift function having an extra parameter $M$ Alfaro et al. (2024). In these two scenarios, using the QNMs obtained, we then verify if we have a parameter-controlled redshift, similar to the environment-induced redshift mentioned above. The modelling of such red-shifts is crucial as the red-shifted modes may appear to be QNMs of the original isolated geometry with a larger throat radius (as $\omega$ varies inversely with $b_{0}$).
Let us first fix our general idea using some well-studied potentials as toy examples. We will try to find out how the $\omega_{QNM}$ vary if we change the width or other parameters present in two well-known potentials which have been investigated and used extensively in the past.
VI.1 Rectangular Barrier:
The rectangular barrier has the form:
$V(x)=\begin{cases}V_{0}&\text{for }0\leq x\leq a,\\ 0&\text{otherwise}.\end{cases}$ (30)
The exact transmission amplitude is $a_{t}=4\omega ke^{-i\omega a}((\omega+k)^{2}e^{-ika}-(\omega-k)^{2}e^{ika})^{-1}$, where $k=\sqrt{\omega^{2}-V_{0}}$.
In order to obtain the quasi-normal spectra, we need purely outgoing waves, i.e. $a_{t}\rightarrow\infty$, Boonserm and Visser (2011) which implies $(\omega+k)^{2}e^{-ika}-(\omega-k)^{2}e^{ika}=0$, and we can solve this equation numerically to get the QNMs of the rectangular barrier.
Here, we are interested in how the QNM spectra change if we change the width (a) of the potential while keeping the height of the barrier fixed (say $V_{0}=1$).
Figure 30: Rectangular Barriers with the same height ($V_{0}=1$) and slightly different widths.
Figure 31: QNMs for Rect. Barrier of different widths. Results of the semi-analytical method and Prony fit are represented by a circle and a plus, respectively.
We have computed the QNM spectra (Fig. 31) for a rectangular barrier with three different barrier widths (Fig. 31). We observe that, as we increase the width of the barrier, the $Re(\omega)$ decreases. Thus, we encounter a redshift for wider barriers. The QNM separation is more pronounced for higher overtones (see Figure 31, rightwards).
VI.2 Poschl-Teller Potential:
The potential is defined as: $V=V_{0}sech^{2}(\alpha x)$ and the corresponding master equation is:
$\frac{d^{2}\psi}{dx^{2}}+\left(\omega^{2}-\frac{V_{0}}{\cosh^{2}(\alpha x)}\right)\psi=0.$
Using $\xi=\tanh(\alpha x)$ and rewriting the equation, we get
$\frac{d}{d\xi}\left((1-\xi^{2})\frac{d\psi}{d\xi}\right)+\left[s(s+1)+\frac{\omega^{2}}{\alpha^{2}(1-\xi^{2})}\right]\psi=0\hskip 28.45274ptwhere\hskip 8.5359pts(s+1)=-\frac{V_{0}}{\alpha^{2}}$ (31)
The asymptotic solution at $x\to-\infty$ is given by Landau and Lifshitz (2013):
$\psi\sim e^{-i\omega x}\frac{\Gamma(\frac{i\omega}{\alpha})\Gamma(1-\frac{i\omega}{\alpha})}{\Gamma(-s)\Gamma(1+s)}+e^{i\omega x}\frac{\Gamma(-\frac{i\omega}{\alpha})\Gamma(1-\frac{i\omega}{\alpha})}{\Gamma(-\frac{i\omega}{\alpha}-s)\Gamma(-\frac{i\omega}{\alpha}+s+1)}$
At the asymptotic limits, the solution must be purely outgoing. Thus, we have to find the values of $\omega$ such that the coefficient of $e^{i\omega x}$ becomes zero. We can solve this analytically, and the QNMs are given by Ferrari and Mashhoon (1984); Beyer (1999); Cardona and Molina (2017)
$\omega=\alpha\left[-i\left(n+\frac{1}{2}\right)\pm\sqrt{\frac{V_{0}}{\alpha^{2}}-\frac{1}{4}}\right]$ (32)
where, $n=0$ corresponds to the fundamental mode and $n=1,2,3,...$ represent overtones.
Figure 32: Poschl-Teller potential of different values of $\alpha$.
Figure 33: QNMs for Poschl-Teller of different values of $\alpha$.
From the plots above, one can note that contrary to the rectangular barrier case, for the Poschl-Teller potential, as we increase the value of $\alpha$ (decrease width), keeping the height fixed (Fig. 33), the real part of the $\omega_{QNM}$ decreases (Fig. 33). So, increasing the value of $\alpha$ in the Poschl-Teller potential causes a redshift.
In both the rectangular barrier and the Poschl-Teller potential case, reducing the height of the potential barrier (keeping the width fixed) causes a reduction in the $Re(\omega)$, thus causing a redshift.
VI.3 Bronnikov–Ellis geometry with asymptotic mass:
@@
Let us now see if a similar redshift effect arises for the QNMs of perturbations of Bronnikov–Ellis geometries. The ultrastatic Bronnikov–Ellis geometry is not of use since it has only one parameter $b_{0}$ (throat radius) which essentially appears as an overall scale factor in the QNMs. Thus, we either modify the shape function with extra parameters, in which case it is no longer Bronnikov–Ellis or, alternatively, we may bring in the redshift function which will have an extra parameter. We choose to work with this latter modification.
@@
where, $f(l)=\frac{C}{b_{0}}\left(arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$. Note that due to the presence of $f(l)$, the energy density as well as the pressures change, as discussed in Section II.2.
@@
The effective potential (scalar perturbation) for this metric can be expressed as Eq. (17). Here, we have plotted the effective potentials for the $n=2$ case ($m=5$ in Fig. 35 and $m=7$ in Fig. 35) and obtained single-peaked potentials with the maxima of the potential shifting to the right as we increase the value of $C$ ($=2M$). The QNMs corresponding to $m=5$ and $m=7$ are plotted in Figs. 37 and 37 respectively.
@@
3 0 $3.53443-0.50686i$ $3.27166-0.46911i$ $3.03576-0.43526i$ $2.82383-0.40465i$ $2.63255-0.37783i$
5 0 $5.52282-0.50282i$ $5.11243-0.46532i$ $4.74378-0.43185i$ $4.41259-0.40161i$ $4.11427-0.37699i$
1 $5.41796-1.52258i$ $5.04381-1.38243i$ $4.65962-1.30864i$ $4.33811-1.21722i$ $4.08699-1.09940i$
2 $5.25123-2.54327i$ $4.85855-2.36301i$ $4.44292-2.19880i$ $4.14809-2.04224i$-
7 0 $7.51934-0.50106i$ $6.95985-0.46386i$ $6.45810-0.43037i$ $6.00711-0.40048i$ $5.60108-0.37349i$
1 $7.45576-1.50817i$ $6.91398-1.38873i$ $6.40556-1.30364i$ $5.95396-1.20411i$ $5.55057-1.11919i$
2 $7.31985-2.52412i$ $6.79114-2.33889i$ $6.28485-2.12915i$ $5.84509-2.02687i$ $5.46827-1.89265i$
9 0 $9.51561-0.50062i$ $8.81008-0.46296i$ $8.17473-0.42973i$ $7.60378-0.39983i$ $7.09023-0.37301i$
1 $9.43351-1.50092i$ $8.76270-1.41073i$ $8.13121-1.29187i$ $7.55941-1.20308i$ $7.04588-1.12225i$
2 $9.35764-2.51354i$ $8.64660-2.31498i$ $8.02958-2.14977i$ $7.50736-2.00358i$ $6.98478-1.87115i$
Table 4: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$. (The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al. (2018b). We extended the analysis to include overtones.]
The effective potential for axial perturbation has the form Eq. (19) and are plotted below. With increasing M ($=\frac{C}{2}$), the height of the barrier reduces and the potentials become more asymmetric (See Figs. 39, 39). The figures 41 and 41 display the QNM spectra for axial perturbation in $m=5$ and $m=7$ cases.
@@
5 0 $5.15176-0.44366i$ $4.77236-0.43237i$ $4.42882-0.40144i$ $4.12030-0.37367i$ $3.84389-0.34924i$
1 $5.06686-1.39798i$ $4.73155-1.33422i$ $4.36034-1.21916i$ $4.06004-1.12686i$ $3.76455-0.99294i$
2 $4.91958-2.39447i$ $4.55509-2.21569i$- -- --
7 0 $7.25073-0.48257i$ $6.71107-0.44682i$ $6.22750-0.41477i$ $5.79355-0.38607i$ $5.40297-0.36046i$
1 $7.19387-1.45787i$ $6.68408-1.34773i$ $6.17347-1.25075i$ $5.74208-1.16371i$ $5.35677-1.08268i$
2 $7.05993-2.43004i$ $6.52963-2.26610i$ $6.04865-2.09636i$ $5.63629-1.94025i$ $5.24520-1.82708i$
9 0 $9.30408-0.48927i$ $8.61413-0.45250i$ $7.99283-0.41976i$ $7.43555-0.39076i$ $6.97356-0.37212i$
1 $9.25255-1.47571i$ $8.56364-1.36779i$ $7.88056-1.26327i$ $7.40840-1.17056i$ $6.88158-1.09045i$
2 $9.15068-2.45648i$ $8.47519-2.27130i$ $7.84263-2.09637i$ $7.27180-1.96059i$ $6.88756-1.86712i$
Table 5: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$. (The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al. (2018b). We extended the analysis to include overtones.]
For this geometry, the increase of asymptotic mass (M) can be associated with redshift, i.e. for higher values of M, we have lower values of $Re(\omega)$. This phenomenon is similar to that observed by the authors of Pezzella et al. (2025). We have presented the quasi-normal modes for small values of M, as we could not get overtones for higher values of M. The fundamental modes are in agreement with the results of Blázquez-Salcedo et al. (2018b).
@@
In Figs. 43 (scalar) and 43 (axial), we have plotted the QNMs (obtained using the Prony Fit method) and the above empirical fitting function for the $n=2$ and $m=7$ case.

### v2

@@ III Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
While computing the $\omega_{QNM}$, we have used a frequency-domain method (spectral method (SM)) in conjunction with a time-domain method (Prony fit) (see appendix A). While the frequency-domain method is indispensable for extracting all the modes (up to a certain overtone number), the time-domain provides us with some valuable insight into the relative amplitudes of the overtones, thus pointing to the fact that sometimes even higher-order overtones (in our case: even modes) can have greater importance than the lower-order (in our case: odd modes) ones.
@@
In the series of figures below, we demonstrate the effective potentials and QNMs for scalar and axial gravitational perturbations. Figs. 3-6 are for $n=2$, Figs. 7-10 for $n=4$, Figs. 11-14 are for $n=6$ and Figs. 15-18 for $n=8$. In the following section, we discuss the results here and try to arrive at specific conclusions on the nature of the QNMs and their link with the different geometries.
@@
IV Inferences drawn from the QNM spectra: emphasis on overtones
[…]
@@ IV Inferences drawn from the QNM spectra: emphasis on overtones
[…]
@@
[…]
@@
[…]
Fixing the value of $n$, if we increase the angular momentum parameter ($m$), the height of the effective potential increases, which is related to increasing the real part of the QNMs. The double peak feature also becomes less prominent for the scalar perturbation case (See Figs. 8, 12, 16). This may explain the decrease of $|Im(\omega)|$ with $m$ in the quasi-normal spectra. For axial perturbations in $n>2$ geometries, the transition of the potential from triple barrier to step-like and then to a single barrier, with increasing m (See Figs. 8, 12, 16), may cause the initial increase and subsequent decrease of $|Im(\omega)|$ for the fundamental modes with m.
IV.1 Approximate Analytic Fit
@@
[…]
@@
Here, we observe that the real parts of the quasi-normal frequencies decrease with overtone mode $p$ but increase with angular momentum parameter $m$ (Fig. 19).
@@
[…]
@@
[…]
IV.2 Detectable Modes
@@
$\nu=\frac{c^{3}}{2\pi b_{0}GM_{\odot}}Re(\omega)\hskip 5.69046ptHz\hskip 56.9055pt\tau=\frac{2\pi b_{0}GM_{\odot}}{c^{3}}\frac{1}{Im(\omega)}\hskip 5.69046ptseconds$
Advanced LIGO is sensitive to the frequency range of $10\hskip 2.84544ptHz$ to $10\hskip 2.84544ptkHz$. For the generalised Bronnikov-Ellis geometries, we have calculated the QNFs, and all of them can be within the aLIGO frequency range when the throat radius is roughly between $50M_{\odot}$ and $1000M_{\odot}$. For now, we choose $b_{0}=200$ (in units of solar mass) and present the frequencies and damping times for a few cases, as an illustration. In Table 2 below, we show the values for $n=2$ and $m=5$ (scalar perturbation).
@@
Table 2: Frequency and Damping time for $n=2$ and $m=5$. These values are calculated from the Prony fit (details in Appendix B).
@@
Frequencies and damping times for n = 8 and m = 5 case (scalar perturbation) are mentioned in Table 3
@@
Table 3: Frequency and Damping time for $n=8$ and $m=5$. These values are calculated from the prony fit (details mentioned in Appendix B).
@@
In the ringdown waveforms, we have represented the actual time-domain (obtained using the procedures mentioned in A.1) wave as the blue points, the fundamental mode as green lines and the resultant wave of the fundamental and overtones ($\omega$ and amplitudes computed using the Prony fit) as the magenta lines (Figs. 23,25). For both cases, the magenta lines better fit the time domain than only the fundamental mode. However, the difference between the green and magenta lines is only noticeable for the first few cycles, as the overtones have higher imaginary parts and decay faster. While the inclusion of overtones makes little improvement in the fitting for the $n=2$ case, the effect is more significant for the $n=8$ case. This is because for the $n=8$ case, we could find more overtones, and these overtones have amplitudes comparable to the fundamental mode, whereas, in the $n=2$ case, the amplitudes of the overtones are significantly lower than the fundamental mode (here the fundamental mode fits well from the second cycle) . For $n=4,6$ cases, we computed 232-3 overtones and their corresponding amplitudes using the Prony fit, and similar to the $n=8$ case, we observe that inclusion of the overtones significantly improves the fit at early times. For all the geometries (i.e. different $n$), the more overtones we could get (up to $p=4$), the better we were able to model the time-domain signal, indicating that obtaining more overtones may result in a better fit of the time-domain profile. However, we need to be cautious not to overfit our results by incorporating an excessive number of overtones.
@@
One would expect the amplitudes of the modes to depend on our initial value of $\psi$ but not on which part of the time domain we are fitting. However, for the generalised Bronnikov–Ellis wormhole, only the amplitudes associated with the fundamental modes and 2nd overtones (to some extent) are time-independent (Figs. 27,27). For almost all the other $\omega_{QNM}$, if we change the initial time of the fitting, the corresponding amplitudes remain stable up to some fitting domain, but beyond it, they begin to vary. Thus, we again fit the ringdown profiles with the modes with a stable amplitude
@@ V Redshift
In the previous sections, we have computed the quasinormal modes for the massless generalised Bronnikov–Ellis wormhole. However, if the geometry under study is surrounded by a dark matter halo or any generic matter distribution, the QNMs calculated for the isolated compact object may differ from the observed values, as the environment may affect the generation of GW as well as their propagation. Such environments cause the spectrum to be red-shifted as documented in Barausse et al. (2014); Pezzella et al. (2025); Liu et al. (2025); Chen and Kotlařík (2023) for the black hole spacetime embedded in a halo.
@@
In the case of the Bronnikov–Ellis wormhole, similar redshifts in the fundamental mode can be found, if one considers the wormhole with non-zero ADM mass, as noted by the authors of Blázquez-Salcedo et al. (2018a). We will first find the QNMs (fundamental and overtones) for the $M\neq 0$ case in the $n=2$ Bronnikov–Ellis geometry. Thereafter, in our second example, we will find the QNMs for a modified $n=2$ geometry with a new redshift function having an extra parameter $M$ Alfaro et al. (2024). In these two scenarios, using the QNMs obtained, we then verify if we have a parameter-controlled redshift, similar to the environment-induced redshift mentioned above. The modelling of such red-shifts is crucial as the red-shifted modes may appear to be QNMs of the original isolated geometry with a larger throat radius (as $\omega$ varies inversely with $b_{0}$).
V.1 Bronnikov–Ellis geometry with asymptotic mass:
@@
Let us now see if redshift effects arise for the QNMs of perturbations of Bronnikov–Ellis geometries. The ultrastatic Bronnikov–Ellis geometry is not of use since it has only one parameter $b_{0}$ (throat radius) which essentially appears as an overall scale factor in the QNMs. Thus, we either modify the shape function with extra parameters, in which case it is no longer Bronnikov–Ellis or, alternatively, we may bring in the redshift function which will have an extra parameter. We choose to work with this latter modification.
@@
where, $f(l)=\frac{C}{b_{0}}\left(\arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$. Note that due to the presence of $f(l)$, the energy density as well as the pressures change, as discussed in Section II.2.
@@
The effective potential (scalar perturbation) for this metric can be expressed as Eq. (17). Here, we have plotted the effective potentials for the $n=2$ case ($m=5$ in Fig. 31 and $m=7$ in Fig. 31) and obtained single-peaked potentials with the maxima of the potential shifting to the right as we increase the value of $C$ (=2M=2M). The QNMs corresponding to $m=5$ and $m=7$ are plotted in Figs. 33 and 33 respectively.
@@
3 0 3.534430.50686i3.53443-0.50686i 3.271660.46911i3.27166-0.46911i 3.035760.43526i3.03576-0.43526i 2.823830.40465i2.82383-0.40465i 2.632550.37783i2.63255-0.37783i
5 0 5.522820.50282i5.52282-0.50282i 5.112430.46532i5.11243-0.46532i 4.743780.43185i4.74378-0.43185i 4.412590.40161i4.41259-0.40161i 4.114270.37699i4.11427-0.37699i
1 5.417961.52258i5.41796-1.52258i 5.043811.38243i5.04381-1.38243i 4.659621.30864i4.65962-1.30864i 4.338111.21722i4.33811-1.21722i 4.086991.09940i4.08699-1.09940i
2 5.251232.54327i5.25123-2.54327i 4.858552.36301i4.85855-2.36301i 4.442922.19880i4.44292-2.19880i 4.148092.04224i4.14809-2.04224i -
7 0 7.519340.50106i7.51934-0.50106i 6.959850.46386i6.95985-0.46386i 6.458100.43037i6.45810-0.43037i 6.007110.40048i6.00711-0.40048i 5.601080.37349i5.60108-0.37349i
1 7.455761.50817i7.45576-1.50817i 6.913981.38873i6.91398-1.38873i 6.405561.30364i6.40556-1.30364i 5.953961.20411i5.95396-1.20411i 5.550571.11919i5.55057-1.11919i
2 7.319852.52412i7.31985-2.52412i 6.791142.33889i6.79114-2.33889i 6.284852.12915i6.28485-2.12915i 5.845092.02687i5.84509-2.02687i 5.468271.89265i5.46827-1.89265i
9 0 9.515610.50062i9.51561-0.50062i 8.810080.46296i8.81008-0.46296i 8.174730.42973i8.17473-0.42973i 7.603780.39983i7.60378-0.39983i 7.090230.37301i7.09023-0.37301i
1 9.433511.50092i9.43351-1.50092i 8.762701.41073i8.76270-1.41073i 8.131211.29187i8.13121-1.29187i 7.559411.20308i7.55941-1.20308i 7.045881.12225i7.04588-1.12225i
2 9.357642.51354i9.35764-2.51354i 8.646602.31498i8.64660-2.31498i 8.029582.14977i8.02958-2.14977i 7.507362.00358i7.50736-2.00358i 6.984781.87115i6.98478-1.87115i
Table 4: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Scalar perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$. (The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al. (2018a). We extended the analysis to include overtones.]
The effective potential for axial perturbation has the form Eq. (19) and are plotted below. With increasing M ($\frac{C}{2}$), the height of the barrier reduces and the potentials become more asymmetric (See Figs. 35, 35). The figures 37 and 37 display the QNM spectra for axial perturbation in $m=5$ and $m=7$ cases.
@@
5 0 5.151760.44366i5.15176-0.44366i 4.772360.43237i4.77236-0.43237i 4.428820.40144i4.42882-0.40144i 4.120300.37367i4.12030-0.37367i 3.843890.34924i3.84389-0.34924i
1 5.066861.39798i5.06686-1.39798i 4.731551.33422i4.73155-1.33422i 4.360341.21916i4.36034-1.21916i 4.060041.12686i4.06004-1.12686i 3.764550.99294i3.76455-0.99294i
2 4.919582.39447i4.91958-2.39447i 4.555092.21569i4.55509-2.21569i - - -
7 0 7.250730.48257i7.25073-0.48257i 6.711070.44682i6.71107-0.44682i 6.227500.41477i6.22750-0.41477i 5.793550.38607i5.79355-0.38607i 5.402970.36046i5.40297-0.36046i
1 7.193871.45787i7.19387-1.45787i 6.684081.34773i6.68408-1.34773i 6.173471.25075i6.17347-1.25075i 5.742081.16371i5.74208-1.16371i 5.356771.08268i5.35677-1.08268i
2 7.059932.43004i7.05993-2.43004i 6.529632.26610i6.52963-2.26610i 6.048652.09636i6.04865-2.09636i 5.636291.94025i5.63629-1.94025i 5.245201.82708i5.24520-1.82708i
9 0 9.304080.48927i9.30408-0.48927i 8.614130.45250i8.61413-0.45250i 7.992830.41976i7.99283-0.41976i 7.435550.39076i7.43555-0.39076i 6.973560.37212i6.97356-0.37212i
1 9.252551.47571i9.25255-1.47571i 8.563641.36779i8.56364-1.36779i 7.880561.26327i7.88056-1.26327i 7.408401.17056i7.40840-1.17056i 6.881581.09045i6.88158-1.09045i
2 9.150682.45648i9.15068-2.45648i 8.475192.27130i8.47519-2.27130i 7.842632.09637i7.84263-2.09637i 7.271801.96059i7.27180-1.96059i 6.887561.86712i6.88756-1.86712i
Table 5: $\omega_{\text{QNM}}$ values for fundamental and overtones for Bronnikov–Ellis wormhole (Axial perturbation) with $M=0,0.025b_{0},0.05b_{0},0.075b_{0},0.1b_{0}$. (The results are obtained using Prony fit.) [The fundamental modes were computed in Blázquez-Salcedo et al. (2018a). We extended the analysis to include overtones.]
@@
In Figs. 39 (scalar) and 39 (axial), we have plotted the QNMs (obtained using the Prony Fit method) and the above empirical fitting function for the $n=2$ and $m=7$ case.